OpenStudy (anonymous):

Chapter 5:Indices and Logarithms Solve the following equation. @ganeshie8

OpenStudy (anonymous):

\[\log _{10}50+2\log _{10}x-\log _{10}(x+4)=2\]

OpenStudy (adamaero):

aren't you forgetting the other two? square the numerator x--right? in the two steps before

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[\log _{10}(\frac{ 50(x^2) }{ x+4 })=2\]

OpenStudy (anonymous):

\[\frac{ 50x^2 }{ x+4 }=100\]

OpenStudy (anonymous):

\[50x^2=400x+400\]

OpenStudy (anonymous):

\[50x^2-400x-400=0\]

OpenStudy (anonymous):

\[50(x^2-8x-8)=0\]

OpenStudy (anonymous):

\[50(x-8)(x+1)=0\]

OpenStudy (anonymous):

\[x=8,-1\]

OpenStudy (adamaero):

After this step log[50(x^2)/(x+4)]=2 why did you multiply each side by 50? I would have used e

ganeshie8 (ganeshie8):

Ecellent! but `log` can only digest positive numbers since you have a term \(\log_{10}(x)\) in your original given equation, you need to throw away the solution \(x=-1\)

OpenStudy (anonymous):

so the answer is only x=8? @ganeshie8

ganeshie8 (ganeshie8):

Yep!

OpenStudy (adamaero):

that's what he's saying but can you explain my question on your question?

ganeshie8 (ganeshie8):

@adamaero is right, there is a mistake after below step : \[\frac{ 50x^2 }{ x+4 }=100\]

OpenStudy (anonymous):

Thnx @ganeshie8 and @adamaero

ganeshie8 (ganeshie8):

\[\frac{ 50x^2 }{ x+4 }=100\] cross multiplying gives u \[50x^2=100(x+4)\]

ganeshie8 (ganeshie8):

cancel 50 both sides and work the remaining quadratic

OpenStudy (anonymous):



OpenStudy (anonymous):

x=4 and x=-2

ganeshie8 (ganeshie8):

yes throw away -2 because it makes the expression \(\log_{10}(x)\) unreal..

OpenStudy (anonymous):

Thnx @ganeshie8 :D